NoFly Corporation sells three different models of a mosquito “zapper.” Model A12 sells for $61 and has variable costs of $40. Model B22 sells for $110 and has variable costs of $72. Model C124 sells for $409 and has variable costs of $311. The sales mix of the three models is A12, 57%; B22, 29%; and C124, 14%.

If the company has fixed costs of $264,312, how many units of each model must the company sell in order to break even? (Round Per unit values to 2 decimal palces, e.g. 15.25 and final answers to 0 decimal places, e.g. 5,275.)

Model
A12 Enter a number of units rounded to 0 decimal places
B22 Enter a number of units rounded to 0 decimal places
C124 Enter a number of units rounded to 0 decimal places
Total break-even

Enter the total break-even in units rounded to 0 decimal places

units
Answer #1

Weighted average contribution margin per unit = (21*57%+38*29%+98*14%) = 36.71

Break even unit = Fixed cost/Weighted average contribution margin per unit = 264312/36.71 = 7200

Model
A12 (7200*57%) 4104
B22 (7200*29%) 2088
C124 (7200*14%) 1008
Total 7200